OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular to the line y = three-fifths x minus 1 and passes through the point (-9, 4). A. five-thirds x plus 19 B. three-fifths x plus forty-seven-fifths C. negative five-thirds x minus 11 D. negative five-thirds x plus 19

OpenStudy (anonymous):

please help this is my last one and im desperate medal to best answer

OpenStudy (mathstudent55):

You are given the line \(y = \dfrac{3}{5} x - 1\) This given line has slope \(\dfrac{3}{5} \). Ok so far?

OpenStudy (mathstudent55):

What is the slope of a perpendicular line?

OpenStudy (anonymous):

y = (3/5)x - 1 Slope is 3/5 The slopes of perpendicular lines are negative reciprocals, so (3/5) m = -1 m = -5/3 So the new slope is -5/3. The new line passes through (-9, 4) y-4 = -5/3(x--9) y-4 = -5/3(x+9) y = -5/3x + 19

OpenStudy (anonymous):

Wait a minute sorry I multiplied wrong. It should be y-4 = -5/3x -15 y = -5/3x - 11 So the answer is C not D. Careful with negative signs when distributing! Also thanks for the medal mathstudent!

OpenStudy (mathstudent55):

You can also do it this way. Once you know the slope of the perpendicular, since you need the answer in slope-intercept form, you can use the slope-intercept equation with the slope of the perpendicular, and with the given point plugged in for x and y. Then you solve for b. \(y = mx + b\) We know that for the perpendicular, \(m = - \dfrac{5}{3} \) Use the given point (-9, 4) and the slope to get: \(4 = -\dfrac{5}{3} (-9) + b\) \(4 = 15 + b\) \(b = -11\) Final answer is: \(y = - \dfrac{5}{3} - 11\)